You manage a risky portfolio with an expected rate of return of 18% and a standard deviation of 31%. The T-bill rate is 4%.

Suppose that your client prefers to invest in your fund a proportion y that maximizes the expected return on the complete portfolio subject to the constraint that the complete portfolio’s standard deviation will not exceed 18%.

a. What is the investment proportion, y? (Round your answer to 2 decimal places.)

b. What is the expected rate of return on the complete portfolio? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Risky portfolio

Expected rate of return = 18%

Standard Deviation = 31%

T-Bill rate = 4%

Maximum portfolio's deviation required = 18%

a). So, proportion of risky portfolio y = 18/31 = 58.06%

Proportion of risk free T-bill = 41.94%

b). Expected return of the complete portfolio = .5806*18 + .4194*4 = 12.12%